# Discretion in the Loop: Human Expertise in Algorithm-Assisted College Advising

### v1

@@ 4.1 Model of Expertise in Interventions
Our model generalizes the functional model of human expertise in prediction, presented in Alur et al. [5], in two ways. First, we fully specify the causal structure among variables using DAGs. Second, we move beyond prediction to intervention, accounting for the possibility that the human decision-maker’s choices can influence future outcomes. [^9] Specifically, we replace the prediction variable $\hat{Y}$ [see eq. 5, 5]—which was assumed to have no effect on the predictive target $Y$—by an action variable $A_{t}$. The action $A_{t}$ may have a causal effect on the future outcome $Y_{t+1}$. Also, the action $A_{t}$ may be informed by both an algorithmic signal (represented by $\mathbf{X}_{t}\to A_{t}$) and additional, unobserved contextual knowledge (represented by $\mathbf{U}_{t}\to A_{t}$). [^10] Our model can capture three types of algorithm-assisted decision making of increasing generality. In the special case $A_{t}$ is performing an idle prediction $\hat{Y}$, we can simply stipulate $A_{t}$ does not influence the outcome $Y_{t+1}$ and our model reduces to Alur et al. [5]’s model. But unlike Alur et al. 5, our model can also represent predictive decision making, where predictions are made of an outcome in order to influence that outcome. With $R(X)$ denoting the range of a random variable $X$, we may think of $A$ as determined by a (potentially noisy) decision policy $\pi:R(\hat{Y})\to R(A)$ on the human’s internal predictions $\hat{Y}$. Finally, our model can represent more general algorithm-assisted interventional contexts, in which actions to affect an outcome do not necessarily depend on predictions of that outcome. Collegiate advising is one such example: An advisor may decide to take actions to support students for many reasons that are not all well-modeled as predictions of the outcome they aim to influence.
@@
Finally, the model specifies a joint distribution $\mathbb{P}$ over the exogenous noise variables $\mathbf{\epsilon_{\mathcal{A}}}$, assumed to be independent. Together with the structural equations, this distribution induces unique observational and causal distributions on all model variables. The observational distribution is Markovian with respect to the causal DAG. [^11] That is, every variable is independent of its non-descendants given its parents. As is standard for causal identification from observational data, we assume the model is causally minimal, that is, the distribution is not Markovian with respect to any proper subgraph [57, pp. 107-109].
@@ 5.1 ExpertTest*: A Conditional Independence Test for Interventions
Given a dataset $\mathcal{D}$ of $n$ $(\mathbf{x}_{t}^{i},a_{t}^{i},y_{t+1}^{i})$ observations, instantiations of the random variables ($\mathbf{X}_{t},A_{t},Y_{t+1}$), we look for independence of actions $A_{t}$ and outcomes $Y_{t+1}$ given the available data to an algorithmic predictor, $\mathbf{X}_{t}$:
$H_{0}:A_{t}\perp Y_{t+1}|\mathbf{X}_{t}$
@@ 5.1.1 ExpertTest* Implementation
We start by constructing a sequence of $K+1$ datasets that would be exchangeable under the null hypothesis $H_{0}$ given above, the first of which $\mathcal{D}_{0}=\{(x_{t}^{1},a_{t}^{1},y_{t+1}^{1})\ldots(x_{t}^{n},a_{t}^{n},y_{t+1}^{n})\}$ is the true dataset, and the other $K$ of which are synthetic datasets $\mathcal{D}_{1}\ldots\mathcal{D}_{K}$. To construct those synthetic datasets, we identify $L$ pairs of individuals with (nearly) identical $\mathbf{X}_{t}$ values (“algorithmically-indistinguishable individuals”) in $\mathcal{D}_{0}$. [^20] When individuals in $\mathcal{D}_{0}$ cannot be paired exactly, as is the case for our dataset, the test can be relaxed to accommodate approximate pairing, introducing error that scales in match quality (measured by a distance function on $\mathbf{X}_{t}$). We refer readers to [5] for the justification of the relaxation as well as rigorous statistical justification of the test. Then, using those $L$ pairs, $K$ synthetic datasets are constructed from $\mathcal{D}_{0}$ by making swaps of the actions advisors take for paired individuals with probability $\frac{1}{2}$.
The mean squared error of a dataset is defined as $MSE(\mathcal{D})=\sum_{i=1}^{n}(a_{t}^{i}-y_{t+1}^{i})^{2}$ with $n=|\mathcal{D}|$. Let $11[\alpha\gtrsim\beta]=$ 1 if $\alpha>\beta$ with ties broken randomly. Given exchangeability, the order statistics over the mean squared error $MSE(\mathcal{D})$ of these datasets,
@@
As discussed in §4.3.3, actions intended to assist at-risk students may be negatively associated with outcomes. This is a key point of departure from the pure prediction setting [5], where good predictions must be positively associated with the outcomes. To address this, for every action $A^{i+}_{t}$ (e.g., recommending a schedule), ExpertTest* is also applied to the transformed action variable $A_{t}^{i-}=1-A_{t}^{i+}$ (e.g., not recommending a schedule). We define
$MSE^{-}(\mathcal{D})=\sum_{i=1}^{n}\left((1-a_{t}^{i})-y_{t+1}^{i}\right)^{2}$
and run the same procedure described above using $MSE^{-}$, allowing us to detect cases where taking an action is associated with worse outcomes, consistent with effective targeting of hidden risks. [^21] Another reasonable choice would be to run a two-tailed test on the order statistic $\tau_{K}$, which we opt against primarily for reasons of interpretability. See Appendix A.3.2 for a comparison; the methods yield similar results.
@@ 5.1.2 Application to MAAPS Data
The test requires matching students with approximately identical algorithmically-available features. We pair students greedily using a Euclidean distance function applied to fixed-range scaled features (Figure 7). [^22] Unlike the procedure for regressions (§7), we do not convert SAT scores to ACT scores, with the effect that students who took different college entry tests are not matched. This allows us to use students’ scores broken down by subject, for which there is no official conversion across tests. We select the number of pairs for each test, $L$, to yield an acceptable tradeoff between test power, which increases in $L$, and test error, which increases in match distance (Appendix A.3.3).
@@ 5.2.3 Discussion: Evidence of Expertise
In the case of a significant result (actions predictive of outcomes in either direction), is there evidence that the interventions are non-algorithmically targeted? While many of the significant interventions present opportunities for advisors to incorporate non-algorithmic information into their decision to apply them, “Other” interventions (Intervention 20), which exhibit significant dependence with term GPA, make for an especially plausible instance of expert targeting. Since the list of interventions is already quite comprehensive, we hypothesize the decision to take “Other” actions might be tailored to bespoke student needs, identified through non-algorithmic information.
@@
In the case the intervention is predictive of negative outcomes, is there evidence that advisors are targeting to at-risk students? Most of the significant interventions predict negative, rather than positive, outcomes. If advisors are not targeting on risk factors (and there is no other hidden confounding), then it is possible these actions of advisors hurt student outcomes. Since the interventions that negatively predict outcomes often signal that the student might be in trouble (like being advised about degree progression issues or advised to come in for an appointment), we hypothesize that advisors might be targeting based on algorithmically-invisible risk factors.
@@
Could any significant results be driven by nuisance confounding? The association of Intervention 17 (being advised about recommended schedule) with persistence suggests some interventions might be confounded by nuisance factors. It may be that students who are likely to persist to next semester are also likely to ask their advisors for schedule advice in the first place, introducing dependence between schedule recommendations and outcomes that is not due to targeting.
@@ 6 Qualitative Analysis: Evidence for Non-Algorithmic Targeted Action
We conduct a two-stage qualitative analysis to better understand advisor discretion and intervention practices, complementing our quantitative findings. The first stage involves open coding to identify themes across all advisor comments, to familiarize ourselves with the data and its limitations. The second stage focuses on a targeted analysis of comments linked to statistically significant interventions from the ExpertTest* to investigate potential evidence of expert targeting. In this section, we describe the methods for our qualitative analysis in §6.1 and how they complement our quantitative findings about advisor expertise in §6.2. In §7.1.2, we discuss results from the exploratory qualitative analysis in connection to our analysis of advisor heterogeneity.
@@ 6.1.1 Exploratory Coding
In the first stage of our qualitative analysis, we conduct exploratory coding on all advisor meetings comments to understand the nature of advisor-student interactions as documented by those comments. Following removal of students due to our data cleaning (§3.1.3), we code a total of 1,594 comments left by advisors on meetings with students in the treatment arm. Using the Atlas.ti software, we apply both deductive codes based on our research questions and inductive codes that emerged from the data. To determine the codes, Authors 1, 2, and 6 first coded subsamples of comments independently, and then met to design a codebook, coordinating to add any additional codes as necessary while coding the entire dataset. We categorize codes into key thematic areas—academic, nonacademic, and relationship-building—further refining these with subcategories and attention to whether the advisor take a directive or exploratory approach.
We note that, while these codes provide valuable insight into advisor styles and the content of meetings, they are necessarily limited to what advisors chose to document, and may not fully capture the scope or nuance of each interaction. Details on our coding process are provided in Appendix A.5. Full code definitions and examples are provided in Appendix A.4.1.
The open coding process in the first round of analysis reveals key themes in advisor-student interactions. By examining the distribution of codes across advisors, student racial groups, and intervention types, we are able to explore variation in advising styles, assess consistency in how interventions were documented, and contextualize observed differences in advisor effectiveness and treatment heterogeneity (more in §7.1.2). This analysis also serves as a check on data quality, showing whether advisors’ written comments aligned with the interventions they reported taking.
@@ 6.1.2 Coding for Expert Targeting
Authors 1 and 2 used a spreadsheet to identify whether advisors explicitly mentioned taking the intervention and provided a reason for doing so (Table 15). For comments that met these requirements, these authors then independently coded the reasons advisors gave for comments associated with different interventions, before meeting to merge the codes and standardize their usage. We evaluate whether these codes indicate (A) the presence of non-algorithmic, outcome-relevant information guiding the choice of intervention, (B) evidence of targeting at-risk students, or (C) the presence of nuisance confounders (e.g., student requests), in accordance with the questions arising from our quantitative analysis. Full code definitions and examples are provided in Appendix A.4.1.
Two interventions require additional coding strategies. Recommending a schedule (Intervention 17) was taken as a default action for most students, which, as discussed in §4.1, trivializes the targeted action component of the expertise definition. Thus we additionally code for customization, identifying whether advisors tailored the content of their recommendations based on student-specific context. For "Other" interventions (Intervention 20), we develop a taxonomy of subtypes of interventions, reporting only those applied frequently enough to ensure consistent interpretation. Details of these specific coding strategies are included in Appendix A.5.1.
To reduce confirmation bias in identifying human expertise, we adopt a conservative approach: we interpret advisor reasoning when only clearly stated and apply the same coding procedures to a sample of statistically insignificant interventions for comparison. [^25] We select the two most insignificant interventions on the ExpertTest* for each outcome with greater than 100 instances of application; one of these overlaps for both outcomes, leading to three total interventions.
@@ 6.2 Qualitative Findings on Non-Algorithmic Targeting
We use the qualitative coding described above to answer three important questions about the mechanisms for a significant result in the ExpertTest*, as described in §5.2.3: (A) whether non-algorithmic targeting is present, (B) whether advisors support at-risk students in case of the intevention predicting negative outcomes, and (C) whether there is evidence of nuisance confounding. We include selected excerpts from advisors’ comments for illustration; comments are paraphrased as needed for clarity and to preserve anonymity. [^26] To protect student privacy, all gendered pronouns in comments have been changed to they/them.
@@
Across all significant interventions, we see a consistent pattern: advisors incorporate non-algorithmic context into their decisions about when and how to intervene. For example, in Intervention 2 (advising a correction to a student’s schedule), advisors incorporate complex information, such as students’ course preferences and history, in their decision to suggest a schedule correction. In one case, an advisor recommends a student postpone a difficult course until the summer based on the student’s self-reflection and planning:
@@
While it is difficult to confirm these factors are not algorithmically predictable from basic administrative data or directly influence outcomes, the complexity of the reasoning behind deciding on an intervention and the variety of the reasons provided for those actions make this plausible in many cases.
@@
“Student in pre-nursing major stated that math was giving them trouble. Without being discouraging, I explained the GPA requirements”.
Moreover, for any given intervention, the distribution of kinds of information that can persuade the advisor to act has a long tail: in most cases, there are many different kinds of information that can lead an advisor to take a particular action, making it difficult to improve a targeting algorithm by simple augmentations of the dataset and data collection process. Table 3 demonstrates the diversity of the reasons advisors cite for taking all the significant interventions we analyzed. When advising a schedule correction (Intervention 2), for instance, advisors consider matters including student’s course preferences, performance, financial aid requirements, projected course loads, prior placement exam performance, and credits earned elsewhere (which often have not been recorded in administrative data at the time of their conversation).
@@
While all significant interventions show signs of non-algorithmic targeting, some also show signs of a confounding effect, in which advisors perform interventions after students explicitly request they do. One such case is recommending a schedule (Intervention 17), which positively predicts next-term persistence ($p$=0.039), as students often seek scheduling advice for the next semester:
@@
Such student requests confound the relationship between actions and outcomes with a factor that is arguably not non-algorithmic information. A student requesting schedule recommendations for next semester from their advisor likely also intends to enroll regardless of the advisor’s particular recommendations. Similar confounding also occurs in comments related to schedule corrections (Intervention 2), progression issues (Intervention 4), and being advised to schedule an appointment (Intervention 16), as the counts in row 12 of Table 3 (student-initiated interventions) demonstrate.
@@
Reason Category Int 2 Int 4 Int 5 Int 9 Int 10 Int 16 Category Sum
1 Course Credits and Placement 3 4 5 - - - 12
2 Course Problem 11 6 6 - - 6 29
3 Degree Progress - 8 3 - - - 11
4 Documents - - - 5 - - 5
5 Fin. Registration Problem - - - 52 - 4 56
6 Load Management 18 - 1 - - - 19
7 Major Choice - 8 1 - - 18 27
8 Non-fin. Registration Problem - - - - 3 - 3
9 Personal Circumstances - 2 - 1 - 5 8
10 Low Engagement - - - - - 109 109
11 Scholarship Process 2 - - 6 - 1 9
12 Student-Initiated 3 - - 1 - 8 12
13 Transfer/Dropout - 3 - - - 3 6
14 Unregistered - - - - - 12 12

### v2

@@ 3.1 Model of Expertise in Interventions
Our model generalizes the functional model of human expertise in prediction, presented in Alur et al. [4], in two ways. First, we fully specify the causal structure among variables using DAGs. Second, we move beyond prediction to intervention, accounting for the possibility that the human decision-maker’s choices can influence future outcomes. [^9] Specifically, we replace the prediction variable $\hat{Y}$ [see eq. 5, 4]—which was assumed to have no effect on the predictive target $Y$—by an action variable $A_{t}$. The action $A_{t}$ may have a causal effect on the future outcome $Y_{t+1}$. Also, the action $A_{t}$ may be informed by both an algorithmic signal (represented by $\mathbf{X}_{t}\to A_{t}$) and additional, unobserved contextual knowledge (represented by $\mathbf{U}_{t}\to A_{t}$). [^10] Our model can capture three types of algorithm-assisted decision making of increasing generality. In the special case $A_{t}$ is performing an idle prediction $\hat{Y}$, we can simply stipulate $A_{t}$ does not influence the outcome $Y_{t+1}$ and our model reduces to Alur et al. [4]’s model. But unlike Alur et al. 4, our model can also represent predictive decision making, where predictions are made of an outcome in order to influence that outcome. With $R(X)$ denoting the range of a random variable $X$, we may think of $A$ as determined by a (potentially noisy) decision policy $\pi:R(\hat{Y})\to R(A)$ on the human’s internal predictions $\hat{Y}$. Finally, our model can represent more general algorithm-assisted interventional contexts, in which actions to affect an outcome do not necessarily depend on predictions of that outcome. Collegiate advising is one such example: An advisor may decide to take actions to support students for many reasons that are not all well-modeled as predictions of the outcome they aim to influence.
@@
Finally, the model specifies a joint distribution $\mathbb{P}$ over the exogenous noise variables $\mathbf{\epsilon_{\mathcal{A}}}$, assumed to be independent. Together with the structural equations, this distribution induces unique observational and causal distributions on all model variables. The observational distribution is Markovian with respect to the causal DAG. [^11] That is, every variable is independent of its non-descendants given its parents. As is standard for causal identification from observational data, we assume the model is causally minimal, that is, the distribution is not Markovian with respect to any proper subgraph [53, pp. 107-109].
@@ 4.1 ExpertTest*: A Conditional Independence Test for Interventions
Given a dataset $\mathcal{D}$ of $n$ $(\mathbf{x}_{t}^{j},a_{t}^{i,j},y_{t+1}^{j})$ observations, instantiations of the random variables ($\mathbf{X}_{t},A_{t}^{i},Y_{t+1}$), we look for independence of actions $A_{t}$ and outcomes $Y_{t+1}$ given the available data to an algorithmic predictor, $\mathbf{X}_{t}$:
$H_{0}:A_{t}^{i}\perp Y_{t+1}|\mathbf{X}_{t}$
@@ 4.1.1 ExpertTest* Implementation
We start by constructing a sequence of $K+1$ datasets that would be exchangeable under the null hypothesis $H_{0}$ given above, the first of which $\mathcal{D}_{0}=\{(x_{t}^{1},a_{t}^{i,1},y_{t+1}^{i,1})\ldots(x_{t}^{n},a_{t}^{i,n},y_{t+1}^{i,n})\}$ is the true dataset, and the other $K$ of which are synthetic datasets $\mathcal{D}_{1}\ldots\mathcal{D}_{K}$. To construct those synthetic datasets, we identify $L$ pairs of individuals with (nearly) identical $\mathbf{X}_{t}$ values (“algorithmically-indistinguishable individuals”) in $\mathcal{D}_{0}$. [^20] When individuals in $\mathcal{D}_{0}$ cannot be paired exactly, as is the case for our dataset, the test can be relaxed to accommodate approximate pairing, introducing error that scales in match quality (measured by a distance function on $\mathbf{X}_{t}$). We refer readers to [4] for the justification of the relaxation as well as rigorous statistical justification of the test. Then, using those $L$ pairs, $K$ synthetic datasets are constructed from $\mathcal{D}_{0}$ by making swaps of the actions advisors take for paired individuals with probability $\frac{1}{2}$.
The mean squared error of a dataset is defined as $MSE(\mathcal{D})=\sum_{j=1}^{n}(a_{t}^{i,j}-y_{t+1}^{j})^{2}$ with $n=|\mathcal{D}|$. Let $11[\alpha\gtrsim\beta]=$ 1 if $\alpha>\beta$ with ties broken randomly. Given exchangeability, the order statistics over the mean squared error $MSE(\mathcal{D})$ of these datasets,
@@
As discussed in §3.3.3, actions intended to assist at-risk students may be negatively associated with outcomes. This is a key point of departure from the pure prediction setting [4], where good predictions must be positively associated with the outcomes. To address this, for every action $A^{i+}_{t}$ (e.g., recommending a schedule), ExpertTest* is also applied to the transformed action variable $A_{t}^{i-}=1-A_{t}^{i+}$ (e.g., not recommending a schedule). This is equivalent to defining
$MSE^{-}(\mathcal{D})=\sum_{j=1}^{n}\left((1-a_{t}^{i,j})-y_{t+1}^{j}\right)^{2}$
and running the same procedure described above using $MSE^{-}$, allowing us to detect cases where taking an action is associated with worse outcomes, consistent with effective targeting of hidden risks. [^21] Another reasonable choice would be to run a two-tailed test on the order statistic $\tau_{K}$, which we opt against primarily for reasons of interpretability. See Appendix A.3.2 for a comparison; the methods yield similar results.
@@ 4.1.2 Application to MAAPS Data
The test requires matching students with approximately identical algorithmically-available features. We pair students greedily using a Euclidean distance function applied to fixed-range scaled features (Figure 6).We select the number of pairs for each test, $L$, to yield an acceptable tradeoff between test power, which increases in $L$, and test error, which increases in match distance (Appendix A.3.3).
@@ 4.2.3 Discussion: Evidence of Expertise
In the case of a significant result (actions predictive of outcomes in either direction), is there evidence that the interventions are non-algorithmically targeted? For instance, do advisors use non-algorithmic information about students to customize the content of students’ schedules (Intervention 17)? This seems plausible, since many factors informing schedule planning, like particular student interests and timing contingencies, could be difficult to predict from basic administrative data about them.
@@
In the case the intervention is predictive of negative outcomes, is there evidence that advisors are targeting to at-risk students? Most of the significant interventions predict negative, rather than positive, outcomes. If advisors are not targeting on risk factors (and there is no other hidden confounding), then it is possible these actions of advisors hurt student outcomes. Since the interventions that negatively predict outcomes signal that the student might be in trouble (like being advised about degree progression issues or advised to come in for an appointment), we hypothesize that advisors might be targeting based on algorithmically-invisible risk factors.
@@
Could any significant results be driven by nuisance confounding? The association of being advised about recommended schedule (Intervention 17) with persistence suggests some interventions might be confounded by nuisance factors. It may be that students who are likely to persist to next semester are also likely to ask their advisors for schedule advice in the first place, introducing dependence between schedule recommendations and outcomes that is not due to targeting.
@@ 5 Qualitative Analysis: Evidence for Non-Algorithmic Targeted Action
We conduct a two-stage qualitative analysis to better understand advisor discretion and intervention practices, complementing our quantitative findings. The first stage involves open coding to identify themes across all advisor comments, to familiarize ourselves with the data and its limitations. The second stage focuses on a targeted analysis of comments linked to statistically significant interventions from the ExpertTest* to investigate potential evidence of expert targeting. In this section, we describe the methods for our qualitative analysis in §5.1 and how they complement our quantitative findings about advisor expertise in §5.2.
@@ 5.1.1 Exploratory Coding
In the first stage of our qualitative analysis, we conduct exploratory coding on all advisor meetings comments to understand the nature of advisor-student interactions as documented by those comments. Following removal of students due to our data cleaning (§2.1.3), we code a total of 1,594 comments left by advisors on meetings with students in the treatment arm. Using the Atlas.ti software, we apply both deductive codes based on our research questions and inductive codes that emerged from the data. To determine the codes, Authors 1 and 5 first coded subsamples of comments independently, and then met to design a codebook, coordinating to add any additional codes as necessary while coding the entire dataset. We categorize codes into key thematic areas—academic, nonacademic, and relationship-building—further refining these with subcategories and attention to whether the advisor take a directive or exploratory approach.
We note that, while these codes provide insight into advisor styles and the content of meetings, they are necessarily limited to what advisors chose to document, and may not fully capture the scope or nuance of each interaction. Details on our coding process are provided in Appendix A.5. Full code definitions are provided in Appendix A.4.1. The open coding process in the first round of analysis familiarized us with key themes in advisor-student interactions. This analysis also serves as a check on data quality, showing whether advisors’ written comments aligned with the interventions they reported taking.
@@ 5.1.2 Coding for Expert Targeting
Author 1 used a spreadsheet to identify whether advisors explicitly mentioned taking the intervention and provided a reason for doing so (Table 14). For comments that met these requirements, the reasons advisors gave for comments associated with different interventions were identified separately for each intervention, and then merged into a standardized codebook across interventions. We evaluate whether these codes indicate (A) the presence of non-algorithmic, outcome-relevant information guiding the choice of intervention, (B) evidence of targeting at-risk students, or (C) the presence of nuisance confounders (e.g., student requests), in accordance with the questions arising from our quantitative analysis. Full code definitions and examples are provided in Appendix A.4.1.
Recommending a schedule (Intervention 17) was taken as a default action for most students, which, as discussed in §3.1, trivializes the targeted action component of the expertise definition. Thus we additionally code for customization, identifying whether advisors tailored the content of their recommendations based on student-specific context.
To reduce confirmation bias in identifying human expertise, we adopt a conservative approach: we interpret advisor reasoning only when clearly stated and apply the same coding procedures to a sample of statistically insignificant interventions for comparison. [^24] We select the two most insignificant interventions on the ExpertTest* for each outcome with greater than 100 instances of application; one of these overlaps for both outcomes, leading to three total interventions.
@@ 5.2 Qualitative Findings on Non-Algorithmic Targeting
We use the qualitative coding process described above to answer three important questions about the mechanisms for a significant result in the ExpertTest*, as described in §4.2.3: (A) whether non-algorithmic targeting is present, (B) whether advisors support at-risk students in case of the intevention predicting negative outcomes, and (C) whether there is evidence of nuisance confounding. We include selected excerpts from advisors’ comments for illustration; comments are paraphrased as needed for clarity and to preserve anonymity. [^25] To protect student privacy, all gendered pronouns in comments have been changed to they/them.
@@
Across significant interventions, we see a pattern: advisors incorporate non-algorithmic context into their decisions about when and how to intervene. For example, in Intervention 2 (advising a correction to a student’s schedule), advisors incorporate complex information, such as students’ course preferences and history, in their decision to suggest a schedule correction. In one case, an advisor recommends a student postpone a difficult course until the summer based on the student’s self-reflection and planning:
@@
While it is difficult to confirm these factors are algorithmically unpredictable from basic administrative data or directly influence outcomes, the complexity of the reasoning behind deciding on an intervention and the variety of the reasons provided for those actions make this plausible in many cases.
@@
“Student in pre-nursing major stated that math was giving them trouble. Without being discouraging, I explained the GPA requirements.”
Moreover, within interventions, the distribution of kinds of information that can persuade the advisor to act often has a long tail: in most cases, there are many different kinds of information that can lead an advisor to take a particular action, making it difficult to improve a targeting algorithm by simple augmentations of the dataset and data collection process. Table 3 demonstrates the diversity of the reasons advisors cite for taking all the significant interventions we analyzed. When advising a schedule correction (Intervention 2), for instance, advisors consider matters including student’s course preferences, performance, financial aid requirements, projected course loads, prior placement exam performance, and credits earned elsewhere (which often have not been recorded in administrative data at the time of their conversation).
@@
While there is a pattern of interventions showing signs of non-algorithmic targeting, some also show signs of a confounding effect, in which advisors perform interventions after students explicitly request they do. One such case is recommending a schedule (Intervention 17), which positively predicts next-term persistence ($p$=0.039), as students often seek scheduling advice for the next semester (Appendix A.5.3):
@@
Such student requests confound the relationship between actions and outcomes with a factor that is arguably not non-algorithmic information. A student requesting schedule recommendations for next semester from their advisor likely also intends to enroll regardless of the advisor’s particular recommendations. Similar confounding also occurs in comments related to schedule corrections (Intervention 2), progression issues (Intervention 4), being advised to schedule an appointment (Intervention 16), and developing the degree planner (Intervention 20) as the counts in row 12 of Table 3 (student-initiated interventions) demonstrate.
@@
Reason Category Int 2 Int 4 Int 5 Int 9 Int 10 Int 16 Int 20 Category Sum
1 Course Credits and Placement 3 4 5 - - - 1 12
2 Course Problem 11 6 6 - - 6 2 29
3 Degree Progress - 8 3 - - - 2 11
4 Documents - - - 5 - - - 5
5 Fin. Registration Problem - - - 52 - 4 - 56
6 Load Management 18 - 1 - - - - 19
7 Major Choice - 8 1 - - 18 - 27
8 Non-fin. Registration Problem - - - - 3 - - 3
9 Personal Circumstances - 2 - 1 - 5 1 8
10 Low Engagement - - - - - 109 - 109
11 Scholarship Process 2 - - 6 - 1 - 9
12 Student-Initiated 3 - - 1 - 8 1 12
13 Transfer/Dropout - 3 - - - 3 - 6
14 Unregistered - - - - - 12 - 12